OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 2(n - 3) + 5 = 3(n - 1)

OpenStudy (hoblos):

2(n - 3) + 5 = 3(n - 1) 2n-6 +5 = 3n-3 2n-1 = 3n-3 n=2